# Comparative Study of Recurrent Neural Networks for Virtual Analog Audio Effects Modeling

### v3

@@ 2.1 Virtual Analog Modeling
A sequence-to-sequence-based model, based on LSTMs, was designed and proposed for the case of another optical compressor, such as the TubeTech CL 1B compressor  [11], to help LSTMs’ memory. The network is based on two LSTM layers: one acts as an encoder, and the other acts as a decoder. The first takes a segment of past input samples together with the controls’ parameters, and the decoder takes the current input sample (in case of single output) or samples (in case of multi-sample output). The encoder, on processing the input, computes the internal states. These states are used as initial states for the second LSTM. These states represent a summary of the encoder input, which helps the decoder to infer the output. The sequence-to-sequence-based model is compared to LSTMs and fully connected networks, resulting in significantly better accuracy. The dataset, in this case, was at $48$kHz sampling rate. The authors initially considered two control parameters, threshold and ratio, while attack and release parameters were fixed. In  [50], the latter parameters were included. The model’s encoder was replaced with a convolutional layer, and the model was compared to TCNs. Sequence-to-sequence-based models allow for more responsiveness and interactivity but at the expense of an increasing computational cost.
@@
Other works consider hybrid approaches. In  [3], fully connected neural networks exploit measurements within the circuits. Here, the structure includes $2$ to $3$ layers embedded within a discrete-time state-space system and explored for diode-based guitar distortion circuits and the lowpass filter of the Korg MS-20 synthesizer. The method required measurements from the devices’ circuits, which are not always accessible, and is moving away from black-box models. Continuing towards gray box approaches,  [51] is based on the Differentiable Digital Signal Processing (DDSP) framework and applies this concept to RC lowpass filter. Subsequently, the idea is applied to the parametric tone section of a guitar amplifier and a nonlinear guitar overdrive circuit in [52]. The technique learns the parameters of a circuit model from measurements of real devices using a Wiener-Hammerstein model consisting of two IIR filters and a feedforward artificial neural network. The IIR filters are, in this way, trainable and derived from the analysis of the circuit. This approach is also used in [53]. In this case, a cascade of biquad filters is trained, employing a Fourier transform, to model the BOSS MT-2 pedal distortion pedal. The model includes four tunable parameters. Differentiable Digital Filters have been applied for the case of a diode clipper circuit [54], showing higher accuracy and similar computation time when compared to traditional physics-based models. In [55], the authors proposed to learn ordinary differential equations governing the first-order and the second-order diode clipper using fully connected layers and RNNs. The formulation similarly uses the discrete-time state-space as in [3]. Neural networks have also been investigated in the case of wave digital filter formulations [2]. A tube amplifier circuit is simulated using multiport nonlinearities inside a wave digital filter and fully connected layers to learn features from a Kirchhoff-domain dataset. Similarly, in  [56], authors combine recurrent neural networks (RNNs) with the Kirchhoff nodal analysis to create a grey-box guitar amplifier model. Lastly, an example of gray box modeling in the case of the LA-2A leveling amplifier is [57], where it is modeled starting from the structure of a traditional digital compressor but using fully connected and RNN to predict the compressor’s parameters.
Gray box methods have the advantage of helping the learning process with knowledge, potentially leading to more accurate and less computationally demanding models. On the other hand, internal measurement of circuit components is not always feasible because it is invasive or difficult to achieve. Infusing external knowledge could not be trivial. For example, in the case of an optical compressor, the attack and release time are governed by a lighting system whose behavior is difficult to model.
@@ 3 Methods
This study specifically investigates recurrent neural networks for virtual analog modeling. In particular, we aim to identify which type of recurrent layer or architectural configuration provides accurate modeling of different types of audio effects, if any. These architectures offer advantages due to their ability to capture time dependencies in the data by utilizing internal states rather than solely relying on the input received at each iteration. In addition, we restrict the experimentation to relatively small networks, which present low computational complexity and low input-output, allowing implementation on real-time consumer-grade digital audio systems. We have developed models that exhibit minimal input-output delay and are well-suited for the typical audio processing workflow that operates on data in blocks. The designed architectures primarily consist of a recurrent layer along with a conditioning block. This architectural choice helps us incorporate control parameters. The designed models are applied to different audio effects and finally evaluated using different metrics. The primary objective of this study is to explore the capabilities of the different recurrent models in learning from raw digital audio and effectively modeling audio effects. These insights can be used to understand the goodness and weakness of these models and determine which can better suit each audio effect. The models presented here, having limited parameters, are not expected to lead to perfectly accurate emulation but a good starting point.
After selecting several types of analog audio effects to be included in this comparative study, we recorded the dataset using analog effects within our availability. We built the dataset using software plugins emulating specific analog effects. Datasets for hardware or software effects were collected using a tool that we have developed for this specific purpose   [^1] https://github.com/stefanofasciani/DGMD.
In the rest of this section, we first detail the four architectures used in this comparative study; then we present the the metrics we used for evaluating the trained models and the specific audio effects from which we collected the datasets. Finally, explain the process we used to train and test the models.
@@ 3.1 Architectures
The following architectures, based on different types of recurrent layers, are compared in this study: LSTM (a), ED (b), LRU (c), and S4D (d). The four architectures are illustrated in Figure 1. ED is based on the concept of internal state sharing  [11]. These architectures follow the same design, consisting of a linear, fully connected layer, a recurrent layer, and a conditioning block.
The number of input samples is identical for all architectures and set to $64$. This choice represents a trade-off: on the one hand, there is the amount of past input information used to generate the current output, where more is typically better; on the other hand, there is the resulting input-output latency, which increases with the number of input samples. Our goal is to keep this latency below $1.33$ ms (i.e., $64$ samples at $48$ kH.). The input of all architectures is an array, including the $64$ most recent input sample, used to generate the single current output sample. Consequently, the model must perform predictions at an audio rate, one for each output audio sample. The models map a vector of $64$ sample in one scalar output sample:
$y_{n}=g(\boldsymbol{x_{n}})=g(x_{n},x_{n-1},...,x_{n-62},x_{n-63})$ (1)
where $n$ is the discrete time index, and $g$ is the overall function describing the models. Before being fed into the recurrent layer (LSTM, LRU, or S4D), the input samples array $\boldsymbol{x_{n}}$ is processed by a fully connected layer that performs a linear projection:
@@
Figure 1: The four compared architectures: LSTM, ED, LRU, and S4D. All architectures present the same structure: a fully connected layer, a recurrent layer, and a conditioning layer. For the case of ED, the input vector is equally split into two vectors, $\boldsymbol{x_{e}}$ and $\boldsymbol{x_{d}}$. The first vector is used to feed two convolutional layers, which are used to compute the states $[h,c]$. These states are combined with the internal states of the LSTM layer included in the decoder. Specifically, a sigmoid function is applied to the LSTM’s states, which are then element-wise multiplied to $[h,c]$. The result of this operation will be the new LSTM’s internal states. The input, identical in all four cases, is an array of the last $64$ input samples $[x_{n},...,x_{n-63}]$, which produces the output sample $y_{n}$. The inputs are firstly linearly projected and then fed to the recurrent layer. Finally, the current output sample is generated through the output layer.
@@
where $\boldsymbol{h_{n}}$ is the hidden state vector, $\boldsymbol{f}$ is the forget gate, $\boldsymbol{i}$ the input gate, and $\boldsymbol{o}$ the output gate, and $\boldsymbol{c}$ is the cell state vector. The forget gate determines what information from previous steps should be kept or discarded. The input gate decides which information from the current step is relevant to add. The output gate determines the value of the next hidden state. Additionally, the cell state is updated by multiplying the forget vector and the input vector with new values that the neural network deems relevant. The forget vector and input vector are obtained using the sigmoid function, denoted by $\gamma$, which determines how much information should be kept. Values of 00 in this vector indicate information to be discarded, while a value of $1$ indicates information to be kept. Furthermore, the $\phi$ represents the hyperbolic tangent function, used to keep the value of the cell state between $-1$ and $1$. Finally, $\boldsymbol{W}$ and $\boldsymbol{U}$ are the weights’ matrices, and $\boldsymbol{b}$ the vector bias terms.
@@
where $\boldsymbol{h_{n}}$ is the hidden state at time $n$ and acts as ’memory’ of the network, $\boldsymbol{o_{n}}$ the output vector. LRU keeps the recurrent layer linear but moves the nonlinearity to a subsequent fully connected layer. Additionally, to further enhance stability, the LRU incorporates a complex-valued diagonal recurrent matrix initialization and an exponential parameterization. At the same time, a normalization scheme for the hidden states is used to boost the efficiency and accuracy [7].
@@
The remarkable ability of SSMs to capture long-range dependencies stems from their utilization of a specific state matrix $\boldsymbol{\bar{W_{A}}}$ known as the ”HiPPO matrix” [58]. This matrix is designed to encode all the past input history in $\boldsymbol{h_{n}}$, finding a map from $\boldsymbol{u_{n}}$ to a higher dimensional space $\boldsymbol{h_{n}}$ that represents the compression of the history. The matrix allows the model to be conceptualized as a convolutional model that decomposes an input signal onto an orthogonal system of smooth basis functions. In this way, the state $\boldsymbol{h_{n}}$ encodes the history of the signal. In our study, we employ the diagonal state space model variant, S4D model  [59], which is defined by parameterizing its state matrix as a diagonal matrix, boosting the model’s efficiency. By doing so, the basis kernels have closed-form formulas represented by normalized Legendre polynomials $L_{n}(t)$, resulting in the SSM decomposing the input signal $\boldsymbol{u_{n}}$ onto an infinitely long set of basis functions that are orthogonal with respect to an exponentially decaying measure. This characteristic gives the SSM the ability to model long-range dependencies effectively.
@@
The conditioning block is then applied just before the output layer, which consists of a linear fully connected layer with one unit, which is described in the following subsection. Initial experiments indicate improved performance when the conditioning block is placed after the recurrent layers rather than before. This suggests that it is more beneficial for the networks to use the information given by the control parameters to project the output of the recurrent layer and determine the extent to which this information influences the final output rather than influencing the inference of the recurrent layer based on the control parameters.
@@ 3.1.1 Conditioning Block
The conditioning block consists of the FiLM method and the GLU layer, as can be seen in Figure 3.
Figure 3: The Conditioning block operates as follows: the vector with conditioning values $\boldsymbol{p}$ is first projected using a linear fully connected layer. This projection results in an output vector with twice the length $\boldsymbol{p}$. This output vector is then divided into two halves, labeled $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$. These vectors, $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$, are utilized to perform an affine transformation on the vector $\boldsymbol{o_{n}}$, coming from the recurrent layer. Subsequently, the transformed vector passes through another linear fully connected layer, producing another output vector that is double the length of its input. This output is also split into two halves, denoted as $\boldsymbol{q_{1}}$ and $\boldsymbol{q_{2}}$. To complete the process, the softsign function is applied to $\boldsymbol{q_{2}}$, and the result is multiplied element-wise with $\boldsymbol{q_{1}}$ to produce $\boldsymbol{o^{c}_{n}}$
@@
$\boldsymbol{\dot{o}_{n}}$ $=(\boldsymbol{\theta}\boldsymbol{o_{n}}+\boldsymbol{\gamma}).$ (8)
@@ 3.2 Losses and Metrics
The models in this study take raw audio samples as input and produce raw audio samples as output. The most common losses used in this case include the Mean Squared Error (MSE), the Mean Absolute Error (MAE), and the Error-to-Signal Ratio (ESR). Based on our knowledge and supported by empirical experiments, there is no evidence suggesting that more complex loss functions could benefit the learning process, especially when modeling the transformation applied to a signal by a generic audio effect. While using specific and different loss functions for different effects may potentially yield better results, we decided to carry out this study using a fixed loss function across architectures and audio effects. This allows for a fair comparison between the trained models, as they all attempt to minimize the same loss function. Therefore, we selected the MSE for this comparative study as it helps to capture the dynamic aspects of the output signal more effectively [50].
In addition, our quantitative evaluation is based on various metrics to assess the models from different perspectives. In the following expressions, $y_{n}$ and $\hat{y}_{n}$ represent the target and predicted signal, and the symbol $N$ represents the overall number of samples in the output signal.
The first metric is the root-mean-square energy (RMSE) error, defined as:
$M_{RMSE}=\sqrt{\frac{1}{N}\sum_{n=1}^{N}(|y_{n}|^{2}-|\hat{y}_{n}|^{2})}$ (9)
RMSE error provides insights into the energy deviation between the target and prediction. The RMSE envelope is calculated using a frame length $W$ and hop size that is both equal to the minibatch size.
The losses and metrics presented so far calculate the average difference between the target and prediction, which may hide large errors that occur only during signal transients. Transients, such as sharp attacks or decays, may represent a small fraction of the overall duration of an audio signal. In this case, we may obtain a relatively low error even with all transients poorly modeled. Therefore, we have included a metric that considers the differences between pairs of spectrums computed on consecutive overlapping windows, as used in the library  [^2] https://github.com/magenta/ddsp developed after  [1]:
@@
where the Short-Time Fourier Transform (STFT) is computed with a window size equal to $2048$ and hop size equal to $512$ samples. This metric emphasizes the ability to accurately predict portions of the output sound with rapid changes.
To account for the frequency response, we use the multi-resolution STFT  [1]:
@@
which compares linear and logarithmic spectral distances with varying frequency resolutions using the L1 norm. The error is calculated as the sum of the absolute values of the vectors. This metric quantifies the error between the spectra employing different resolutions. We compute two separate multi-resolution STFT errors: $M_{STFT-1}$ with FFT window lengths of $m=[32,64,128]$, and $M_{STFT-2}$ with $m=[256,512,1024]$. This differentiation allows us to evaluate errors occurring in windows with high frequency but low time resolution separately from errors in windows with low frequency but high time resolution. For both cases, the hop size is $25\%$ of the window length.
Similarly, we compare the MFCCs, to compute a metric linked to human sound perception:

### v5

@@ 2.1 Neural Networks for Virtual Analog Modeling
A sequence-to-sequence-based model, based on LSTMs, was designed and proposed for the case of another optical compressor, such as the TubeTech CL 1B compressor  (Simionato & Fasciani,, 2022), to help LSTMs’ memory. The network is based on two LSTM layers: one acts as an encoder, and the other acts as a decoder. The first takes a segment of past input samples together with the controls’ parameters, and the decoder takes the current input sample (in case of single output) or samples (in case of multi-sample output). The encoder, when processing the input, computes the internal states. These states are used as initial states for the second LSTM. These states represent a summary of the encoder input, which helps the decoder to infer the output. The sequence-to-sequence-based model is compared to LSTMs and fully connected networks, resulting in significantly better accuracy. The authors initially considered two control parameters, threshold, and ratio, while attack and release parameters were fixed. In  (Simionato & Fasciani,, 2023), the latter parameters were included. The model’s encoder was replaced with a convolutional layer, and the model was compared to TCNs. Sequence-to-sequence-based models allow for more responsiveness and interactivity but at the expense of an increasing computational cost.
@@
Other works consider hybrid approaches. In  (Parker et al.,, 2019), fully connected neural networks exploit measurements within the circuits. Here, the network is embedded within a discrete-time state-space system and explored for diode-based guitar distortion circuits and the lowpass filter of the Korg MS-20 synthesizer. The method required measurements from the devices’ circuits, which are not always accessible. Continuing towards gray box approaches,  (Kuznetsov et al.,, 2020) is based on the Differentiable Digital Signal Processing (DDSP) framework and applies this concept to RC lowpass filter. Subsequently, the idea is applied to the parametric tone section of a guitar amplifier and a nonlinear guitar overdrive circuit in (Esqueda et al.,, 2021). The technique learns the parameters of a circuit model from measurements of real devices using a Wiener-Hammerstein model consisting of two IIR filters and a feedforward artificial neural network. The IIR filters are, in this way, trainable and derived from the analysis of the circuit. This approach is also used in (Nercessian et al.,, 2021). In this case, a cascade of biquad filters is trained, employing a Fourier transform, to model the BOSS MT-2 pedal distortion pedal. The model includes four tunable parameters. Differentiable Digital Filters have been applied for the case of a diode clipper circuit (Chowdhury & Clarke,, 2022), showing higher accuracy and similar computation time when compared to traditional physics-based models. In (Wilczek et al.,, 2022), the authors proposed to learn ordinary differential equations governing the first-order and the second-order diode clipper using fully connected layers and RNNs. The formulation similarly uses the discrete-time state-space as in (Parker et al.,, 2019). Neural networks have also been investigated in the case of wave digital filter formulations (Darabundit et al.,, 2022). A tube amplifier circuit is simulated using multiport nonlinearities inside a wave digital filter and fully connected layers to learn features from a Kirchhoff-domain dataset. Similarly, in  (Miklanek et al.,, 2023), authors combine recurrent neural networks (RNNs) with the Kirchhoff nodal analysis to create a grey-box guitar amplifier model. Lastly, an example of gray box modeling in the case of the LA-2A leveling amplifier is (Wright & Välimäki,, 2022), where it is modeled starting from the structure of a traditional digital compressor but using fully connected and RNN to predict the compressor’s parameters.
@@ 3 Methods
This study aims to identify which type of recurrent layer or architectural configuration accurately models different types of audio effects, if any. These insights can be used to understand the goodness and weakness of these models and determine which can better suit each audio effect. The states-based architectures offer advantages due to their ability to capture time dependencies in the data by utilizing internal states rather than solely relying on the input received at each iteration. In addition, we restrict the experimentation to relatively small networks fed by a small number of input segments, which present low computational complexity and low input-output latency, allowing implementation on real-time consumer-grade digital audio systems. The designed models are applied to different audio effects and finally evaluated using different metrics. To be noticed, the models presented here, having limited parameters, are not expected to lead to perfectly accurate emulation but a good starting point.
In the rest of this section, we first detail the four architectures used in this comparative study; then, we present the metrics we used for evaluating the trained models and the specific audio effects from which we collected the datasets. Finally, explain the process we used to train and test the models.
@@ 3.1 Architectures
The following architectures, based on different types of recurrent layers, are compared in this study: LSTM (a), ED (b), LRU (c), and S4D (d). The four architectures are illustrated in Figure 1. ED is based on the concept of internal state sharing  (Simionato & Fasciani,, 2022). These architectures follow the same design, consisting of a linear, fully connected layer, a recurrent layer, and a conditioning block.
The number of input samples is identical for all architectures and set to $64$. This choice represents a trade-off: on the one hand, there is the amount of past input information used to generate the current output, where more is typically better; on the other hand, there is the resulting input-output latency, which increases with the number of input samples. Our goal is to keep this latency below $1.33$ ms (i.e., $64$ samples at $48$ kH.). The input of all architectures is an array, including the $64$ most recent input sample, used to generate the single current output sample. This implies that a new input sample enters the buffer with each inference cycle, while the oldest is discarded. Therefore, the set of the $64$ most recent input samples constitutes the network input, which, after initial compression by the FC layer, is used to update the states in the recurrent layers. This approach aids the network in making predictions while maintaining a stateful design using truncated backpropagation through time. Predicting one sample at a time helps minimize the audible artifacts commonly associated with machine-learning models of audio effects, which arise from slight amplitude mismatches at the boundaries of consecutive output segments. Consequently, the model must perform predictions at an audio rate, one for each output audio sample. The models map a vector of $64$ sample in one scalar output sample, give the internal states $\boldsymbol{h}$:
$y_{n}=g(\boldsymbol{x_{n}})=g(x_{n},x_{n-1},...,x_{n-62},x_{n-63};\boldsymbol{h})$ (1)
where $n$ is the discrete time index, and $g$ is the overall function describing the models. Before being fed into the recurrent layer (LSTM, LRU, S4D, or S6), the input samples array $\boldsymbol{x_{n}}$ is processed by a fully connected layer that performs a linear projection:
@@
Figure 1: The four compared architectures: LSTM, ED, LRU, S4D, and S6. All architectures present the same structure: a fully connected layer, a recurrent layer, and a conditioning layer. For the case of ED, the input vector is equally split into two vectors, $\boldsymbol{x_{e}}$ and $\boldsymbol{x_{d}}$.
@@
where $\boldsymbol{h_{n}}$ is the hidden state vector, $\boldsymbol{f}$ is the forget gate, $\boldsymbol{i}$ the input gate, and $\boldsymbol{o}$ the output gate, and $\boldsymbol{c}$ is the cell state vector. The forget gate determines what information from previous steps should be kept or discarded. The input gate decides which information from the current step is relevant to add. The output gate determines the value of the next hidden state. Additionally, the cell state is updated by multiplying the forget vector and the input vector with new values that the neural network deems relevant. The forget vector and input vector are obtained using the sigmoid function, denoted by $\gamma$, which determines how much information should be kept. Values of 00 in this vector indicate information to be discarded, while a value of $1$ indicates information to be kept. Furthermore, the $\phi$ represents the hyperbolic tangent function, used to keep the value of the cell state between $-1$ and $1$. Finally, $\boldsymbol{W}$ and $\boldsymbol{U}$ are the weights’ matrices, and $\boldsymbol{b}$ the vector bias terms. After the LSTM layer, a linear FC layer reduces the dimensionality down to $4$, ensuring consistency with all other models before proceeding to the conditioning block.
@@
where $\boldsymbol{h_{n}}$ is the hidden state at time $n$ and acts as ’memory’ of the network, $\boldsymbol{o_{n}}$ the output vector. LRU keeps the recurrent layer linear but moves the nonlinearity to a subsequent fully connected layer. Additionally, to further enhance stability, the LRU incorporates a complex-valued diagonal recurrent matrix initialization and an exponential parameterization. At the same time, a normalization scheme for the hidden states is used to boost the efficiency and accuracy (Orvieto et al.,, 2023).
@@
In our study, we employ the diagonal state space model variant S4D  (Gu et al.,, 2022a), and newer S6 variant (Gu & Dao,, 2023).
The remarkable ability of S4D to capture long-range dependencies stems from their utilization of a specific state matrix $\boldsymbol{\bar{W_{A}}}$ known as the "HiPPO matrix" (Gu et al.,, 2020). This matrix is designed to encode all the past input history in $\boldsymbol{h_{n}}$, finding a map from $\boldsymbol{u_{n}}$ to a higher dimensional space $\boldsymbol{h_{n}}$ that represents the compression of the history. The matrix allows the model to be conceptualized as a convolutional model that decomposes an input signal onto an orthogonal system of smooth basis functions. In this way, the state $\boldsymbol{h_{n}}$ encodes the history of the signal. In the diagonal form, the S4D model is defined by parameterizing its state matrix as a diagonal matrix, boosting the model’s efficiency. By doing so, the basis kernels have closed-form formulas represented by normalized Legendre polynomials $L_{n}(t)$, resulting in the SSM decomposing the input signal $\boldsymbol{u_{n}}$ onto an infinitely long set of basis functions that are orthogonal with respect to an exponentially decaying measure.
In contrast, the S6 layer introduces input-dependent matrices. Specifically, $\boldsymbol{\bar{W_{B}}}$ and $\boldsymbol{\bar{W_{C}}}$ become dependent on the input and computed using a linear FC layer, while $\boldsymbol{\bar{W_{A}}}$ and $\boldsymbol{\bar{W_{D}}}$’s parameters are independent of the input and learned during the training process.
@@
The conditioning block is then applied just before the output layer, which consists of a linear fully connected layer with one unit, which is described in the following subsection. Placing the conditioning block after the recurrent layers rather than before has been found beneficial  (Fasciani et al.,, 2024a). This suggests that it is more beneficial for the networks to use the information given by the control parameters to project the output of the recurrent layer and determine the extent to which this information influences the final output rather than influencing the inference of the recurrent layer based on the control parameters.
@@ 3.1.1 Conditioning Block
The conditioning block consists of the FiLM method and the GLU layer, as in (Fasciani et al.,, 2024a), and is illustrated in Figure 3.
Figure 3: Conditioning block based on the FiLM method and the GLU layer. The GLU layer features the softsign activation, $\boldsymbol{p}$ denotes the conditioning vector, and $\boldsymbol{o_{n}}$ represents the network’s internal representation of the input.
@@
$\boldsymbol{\dot{o}_{n}}$ $=(\boldsymbol{\theta}\boldsymbol{o_{n}}+\boldsymbol{\eta}).$ (8)
@@
Lastly, the presented models are designed to have limited computational complexity for the inference process. Specifically, we set the limit to $1.500$ FLOPs per sample or $72$ MFLOPs per second at $48$ KHz. By doing so, the models feature $\sim 750$ internal parameters. Table 1 reports the FLOPs per sample for each model.
Table 1: FLOPs per sample for each model considered in this study.
Model Recurrent Layer Conditioning Block Total
LSTM $1040$ $120$ $1160$
ED $928$ $120$ $1048$
LRU $692$ $120$ $812$
S4D $792$ $120$ $912$
S6 $864$ $120$ $984$
@@ 3.2 Losses and Metrics
The models in this study take raw audio samples as input and produce raw audio samples as output. The most common losses used in this case include the Mean Squared Error (MSE), the Mean Absolute Error (MAE), and the Error-to-Signal Ratio (ESR). Based on our knowledge and supported by empirical experiments, there is no evidence suggesting that more complex loss functions could benefit the learning process, especially when modeling the transformation applied to a signal by a generic audio effect. While using specific and different loss functions for different effects may potentially yield better results, we decided to carry out this study using a fixed loss function across architectures and audio effects. This allows for a fair comparison between the trained models, as they all attempt to minimize the same loss function. Therefore, we selected the MSE for this comparative study as it helps to capture the dynamic aspects of the output signal more effectively  (Simionato & Fasciani,, 2023).
In addition, our quantitative evaluation is based on various metrics to assess the models from different perspectives. In the following expressions, $y_{n}$ and $\hat{y}_{n}$ represent the target and predicted signal and the symbol $N$ represents the overall number of samples in the output signal.
The first metric is the root-mean-square energy (RMSE) error. RMSE error provides insights into the energy deviation between the target and prediction. The losses and metrics presented so far calculate the average difference between the target and prediction, which may hide large errors that occur only during signal transients. Transients, such as sharp attacks or decays, may represent a small fraction of the overall duration of an audio signal. In this case, we may obtain a relatively low error even with all transients poorly modeled. Therefore, we have included a metric related to the spectral flux, which considers the differences between pairs of spectrums computed on consecutive overlapping windows, as used in the library  [^1] https://github.com/magenta/ddsp developed after  (Engel et al.,, 2020):
@@
where the Short-Time Fourier Transform (STFT) is computed with a window size equal to $2048$ and hop size equal to $512$ samples. This metric emphasizes the ability to predict portions of the output sound with rapid changes.
To account for the frequency response, we use the multi-resolution STFT  (Engel et al.,, 2020):
@@
which compares linear and logarithmic spectral distances with varying frequency resolutions using the L1 norm. The error is calculated as the sum of the absolute values of the vectors. This metric quantifies the error between the spectra employing different resolutions. We compute the metrics using $m=[256,512,1024]$ as resolutions.
Similarly, we compare the MFCCs to compute a metric linked to human sound perception: